# General Geospatial Inference with a Population Dynamics Foundation Model

### v5

@@ 6.3.3 Comparison Approach
We selected two geographic encoders that showed strong performance on income, housing prices, elevation, and population density in previous experiments (Klemmer et al. (2023)). We evaluate the same four tasks in our experiments in addition to 23 other tasks as part of our benchmark.
Satellite Contrastive Location-Image Pretraining (SatCLIP) by Klemmer et al. (2023) is a geographic encoder based on satellite imagery. SatCLIP outperforms a previous version called MOSAIKS (Rolf et al. (2021)) along with a technique using contrastive self-supervision (CSP) (Mai et al. (2023)), that was trained on species and satellite imagery.
GeoCLIP (Vivanco Cepeda et al. (2024)) is a geographic encoder trained with 4.72 million Flickr images. GeoCLIP often outperforms SatCLIP on income and housing benchmarks.
@@
Inverse Distance Weighting (IDW) (Shepard (1968)) is a technique that is able to interpolate by learning a weighted average of known points. In our case IDW operates by learning only on the latitude and longitude coordinates along with the downstream task labels within our training and validation splits without any additional data. These results are presented alongside our foundation model results.
Radial Basis Function (RBF) Interpolation (Fasshauer (2007)) uses an RBF centered at a target location in combination with a polynomial function to fill in missing values. These results are not presented in this paper because of significantly lower performance compared with IDW.
Gaussian Process Regression (Chung et al. (2019)) incorporates the spatial correlation of sampled points when predicting missing values. These results are not presented in this paper because of significantly lower performance compared with IDW.
@@ 6.4 Application of PDFM Embeddings
Forecasting. We evaluated the utility of the PDFM embeddings by augmenting an existing forecasting model. We selected TimesFM (Time Series Foundation Model) (Das et al. (2023)). TimesFM is a state-of-the-art pre-trained decoder attention model utilizing patched inputs from a large real-world and synthetic time-series corpus. It required no training when applied to a new forecasting problem (although incorporating our PDFM embeddings does require training a simple adapter model, such as the multilayer perceptron (MLP), described below in our experimental setup). As a benchmark and way to calibrate our results, we also incorporated a traditional supervised statistical forecasting approach, ARIMA (AutoRegressive Integrated Moving Average) (Box and Jenkins (1976)). ARIMA leverages past values of the input data and past forecast errors to capture trends and account for noise within the data.
@@ 7 Data and Code Availability
We are publicly releasing the PDFM embeddings, as well as the processed benchmark data for all tasks. Code samples for using the PDFM embeddings and recreating the benchmarks will also be made available for public use. Starting on November 14, 2024, the data and code can be accessed at https://github.com/google-research/population-dynamics.
@@ 8 Competing Interests
IDW SatCLIP GeoCLIP
Weather
& AQ
Trends Maps PDFM
PDFM w/
SatCLIP
@@
IDW SatCLIP GeoCLIP
Weather
& AQ
Trends Maps PDFM
PDFM w/
SatCLIP
@@
IDW SatCLIP GeoCLIP
Weather
& AQ
Trends Maps PDFM
PDFM w/
SatCLIP

### v6

@@ 6.3.3 Comparison Approach
We selected two geographic encoders that showed strong performance on income, housing prices, elevation, and population density in previous experiments (klemmer2023satclip). We evaluate the same four tasks in our experiments in addition to 23 other tasks as part of our benchmark.
Satellite Contrastive Location-Image Pretraining (SatCLIP) by klemmer2023satclip is a geographic encoder based on satellite imagery. SatCLIP outperforms a previous version called MOSAIKS (rolf2021generalizable) along with a technique using contrastive self-supervision (CSP) (mai2023csp), that was trained on species and satellite imagery.
GeoCLIP (cepeda2024geoclip) is a geographic encoder trained with 4.72 million Flickr images. GeoCLIP often outperforms SatCLIP on income and housing benchmarks.
@@
Inverse Distance Weighting (IDW) (shepard1968two) is a technique that is able to interpolate by learning a weighted average of known points. In our case IDW operates by learning only on the latitude and longitude coordinates along with the downstream task labels within our training and validation splits without any additional data. These results are presented alongside our foundation model results.
Radial Basis Function (RBF) Interpolation (fasshauer2007meshfree) uses an RBF centered at a target location in combination with a polynomial function to fill in missing values. These results are not presented in this paper because of significantly lower performance compared with IDW.
Gaussian Process Regression (chung2019supplement) incorporates the spatial correlation of sampled points when predicting missing values. These results are not presented in this paper because of significantly lower performance compared with IDW.
@@ 6.4 Application of PDFM Embeddings
Forecasting. We evaluated the utility of the PDFM embeddings by augmenting an existing forecasting model. We selected TimesFM (Time Series Foundation Model) (das2023decoder). TimesFM is a state-of-the-art pre-trained decoder attention model utilizing patched inputs from a large real-world and synthetic time-series corpus. It required no training when applied to a new forecasting problem (although incorporating our PDFM embeddings does require training a simple adapter model, such as the multilayer perceptron (MLP), described below in our experimental setup). As a benchmark and way to calibrate our results, we also incorporated a traditional supervised statistical forecasting approach, ARIMA (AutoRegressive Integrated Moving Average) (box1976time). ARIMA leverages past values of the input data and past forecast errors to capture trends and account for noise within the data.
@@ 7 Data and Code Availability
We offer a program for academic and non-commercial researchers to access PDFM embeddings. To learn more, please email pdfm-embeddings@google.com. Once access is granted, public code samples for using the PDFM embeddings and creating the benchmarks can be found at https://github.com/google-research/population-dynamics.
@@ 8 Competing Interests
IDW SatCLIP GeoCLIP Weather & AQ Trends Maps PDFM PDFM w/ SatCLIP
@@
IDW SatCLIP GeoCLIP Weather & AQ Trends Maps PDFM PDFM w/ SatCLIP
@@
IDW SatCLIP GeoCLIP Weather & AQ Trends Maps PDFM PDFM w/ SatCLIP